# Topological Data Analysis in Finsler Spaces

### v1

@@ 2 Finsler Structure
    $F$ is p-homogeneous of degree $1$ in $v$, that is, $F(p,rv)=rF(p,v),\leavevmode\nobreak\ \forall r>0$;
@@
$d_{F}(x,y)=\inf_{\gamma}\leavevmode\nobreak\ L(\gamma),$ (2)
@@
For the Euclidean Finsler metric, the curve $\gamma$ that satisfies the solutions of Euler-Lagrange equations are $\gamma^{i}(t)=a_{i}\leavevmode\nobreak\ t+b_{i}$ for constants $a_{i}$ and $b_{i}$. For two points $x,y\in M$ with $\gamma(0)=x^{i}$ and $\gamma(1)=y^{i}$, we have $a_{i}=y^{i}-x^{i}$ and $b^{i}=x^{i}$, then
$d_{E}(x,y)=\int_{0}^{1}E(a_{i}\leavevmode\nobreak\ t+b_{i},a_{i})\leavevmode\nobreak\ dt=\int_{0}^{1}E(t(y^{i}-x^{i})+x^{i},y^{i}-x^{i})\leavevmode\nobreak\ dt=\left(\sum_{i=1}^{n}(y^{i}-x^{i})^{2}\right)^{\frac{1}{2}},$ (4)
@@ 5 Finsler Structures in TDA
Proof. If $F(p,v)$ does not depend on first variable then the curves that perform the distance $d_{F}$ satisfies $\dfrac{d^{2}p^{i}}{ds^{2}}=0,$ i.e., the curves are $\gamma^{i}(s)=a_{i}\leavevmode\nobreak\ s+b_{i}$ for $a_{i},b_{i}$ constants. If $\gamma^{i}(0)=x^{i}$, $\gamma^{i}(1)=y^{i}$, we have $a_{i}=y^{i}-x^{i}$ and $b^{i}=x^{i}$, then
$d_{F}(x,y)=\int_{0}^{1}F(a_{i}\leavevmode\nobreak\ s+b_{i},a_{i})\leavevmode\nobreak\ ds=\int_{0}^{1}F(s(y^{i}-x^{i})+x^{i},y^{i}-x^{i})\leavevmode\nobreak\ ds$ (12)
$=\int_{0}^{1}F(y^{i}-x^{i})\leavevmode\nobreak\ ds=s\leavevmode\nobreak\ F(y^{i}-x^{i})\Big{|}_{0}^{1}=F(y-x)=F_{x}(y-x).$ (13)
@@ 6 Stability of the persistent diagrams
$d_{b}(dgm(U),dgm(V))=\inf_{\eta}\leavevmode\nobreak\ \sup_{x}\parallel x-\eta(x)\parallel_{\infty},$

### v2

@@ 2 Finsler Structure
    $F$ is p-homogeneous of degree $1$ in $v$, that is, $F(p,rv)=rF(p,v),\penalty 10000\ \forall r>0$;
@@
$d_{F}(x,y)=\inf_{\gamma}\penalty 10000\ L(\gamma),$ (2)
@@
For the Euclidean Finsler metric, the curve $\gamma$ that satisfies the solutions of Euler-Lagrange equations are $\gamma^{i}(t)=a_{i}\penalty 10000\ t+b_{i}$ for constants $a_{i}$ and $b_{i}$. For two points $x,y\in M$ with $\gamma(0)=x^{i}$ and $\gamma(1)=y^{i}$, we have $a_{i}=y^{i}-x^{i}$ and $b^{i}=x^{i}$, then
$d_{E}(x,y)=\int_{0}^{1}E(a_{i}\penalty 10000\ t+b_{i},a_{i})\penalty 10000\ dt=\int_{0}^{1}E(t(y^{i}-x^{i})+x^{i},y^{i}-x^{i})\penalty 10000\ dt=\left(\sum_{i=1}^{n}(y^{i}-x^{i})^{2}\right)^{\frac{1}{2}},$ (4)
@@ 5 Finsler Structures in TDA
Proof. If $F(p,v)$ does not depend on first variable then the curves that perform the distance $d_{F}$ satisfies $\dfrac{d^{2}p^{i}}{ds^{2}}=0,$ i.e., the curves are $\gamma^{i}(s)=a_{i}\penalty 10000\ s+b_{i}$ for $a_{i},b_{i}$ constants. If $\gamma^{i}(0)=x^{i}$, $\gamma^{i}(1)=y^{i}$, we have $a_{i}=y^{i}-x^{i}$ and $b^{i}=x^{i}$, then
$d_{F}(x,y)=\int_{0}^{1}F(a_{i}\penalty 10000\ s+b_{i},a_{i})\penalty 10000\ ds=\int_{0}^{1}F(s(y^{i}-x^{i})+x^{i},y^{i}-x^{i})\penalty 10000\ ds$ (12)
$=\int_{0}^{1}F(y^{i}-x^{i})\penalty 10000\ ds=s\penalty 10000\ F(y^{i}-x^{i})\Big|_{0}^{1}=F(y-x)=F_{x}(y-x).$ (13)
@@ 6 Stability of the persistent diagrams
$d_{b}(dgm(U),dgm(V))=\inf_{\eta}\penalty 10000\ \sup_{x}\parallel x-\eta(x)\parallel_{\infty},$